# Transforming Code Repositories into MCP Services

### v3

@@ Abstract
The Model Context Protocol (MCP) aims to create a standard for how Large Language Models use tools. However, most current research focuses on selecting tools from an existing pool. A more fundamental, yet largely overlooked, problem is how to populate this pool by converting the vast number of existing software projects into MCP-compatible services. To bridge this gap, we introduce Code2MCP, an agent-based framework that automatically transforms a GitHub repository into a functional MCP service with minimal human intervention. Code2MCP employs a multi-agent workflow for code analysis, environment setup, tool function design, and service generation, enhanced by a self-correcting loop to ensure reliability. We demonstrate that Code2MCP successfully transforms open-source computing libraries in scientific fields such as bioinformatics, mathematics, and fluid dynamics that are not available in existing MCP servers. By providing a novel automated pathway to unlock GitHub, the world’s largest code repository, for the MCP ecosystem, Code2MCP serves as a catalyst to significantly accelerate the protocol’s adoption and practical application. The code is public at https://github.com/DEFENSE-SEU/Code2MCP.
1 Introduction
@@
The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024; bubeck2023sparks). To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023). Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022).
@@
However, this burgeoning ecosystem faces a fundamental scalability challenge: the $N\times M$ integration problem (Li et al., 2023; Liang et al., 2023; Qin et al., 2023; Anthropic, 2023). Each of the N models or agent applications often requires a bespoke connector for each of the M tools it must access. This results in a fragmented and inefficient system where development effort is duplicated and innovation is stifled by high integration costs (Qu et al., 2025; Shen, 2024). To address this, MCP is proposed as a universal standard that specifies how agents and tools should communicate, enabling an interoperable “plug-and-play” ecosystem (Anthropic, 2023).
In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025). The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2023; Ding et al., 2025; Qin et al., 2024). To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang and others, 2025; Xie et al., 2023; Wang, 2023). This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code. The MCP emerges as a direct answer to this chaos, promising a universal interface (Yu et al., 2025; Anthropic, 2023). Yet, this leads to the critical gap: research is focused almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational “supply-side” problem of how to populate this pool from existing software is largely unaddressed.
@@
Figure 1: While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side). Code2MCP solves the supply problem by converting the code repository into a standardized MCP-compliant tool.
@@
However, while these efforts advance the consumption side of the problem, how agents can better use tools, they largely overlook a more fundamental bottleneck on the supply side. This supply bottleneck is not a theoretical concern but a stark reality preventing standards like MCP from achieving widespread adoption. For example, RAG-MCP (Gan and Sun, 2025) utilizes over 4,400 servers on mcp.so, but there are 268 million public GitHub repositories. The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed. This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
@@
In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap. Code2MCP presents a blueprint for transforming any GitHub repository into a functional and documented MCP service with minimal human intervention. However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process. To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023). Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024b; Huang et al., 2023), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design. Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, which endows the system with the ability to autonomously debug and refine the entire conversion process. The key contributions of this work are listed as follows:
@@
Table 1: A comparative summary of related works in tool-augmented LLMs.
Work Core Contribution Tool Pool Tool Selection MCP
Consumer Toolformer Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
SciToolAgent Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
HuggingGPT Increase the size of tool pool huggingface.co LLM task planning $\times$
Gorilla TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
OpenAgents Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
RepoMaster Rule-based deep search $\times$
RAG-MCP RAG for tool selection from MCP MCP.so (4,400+) Retrieval on MCP $\checkmark$
2 Related Work
@@
Scaling Tool Availability via Structured Platforms. To overcome the limitation of fixed toolsets, subsequent research turns to large, curated platforms. These approaches significantly expand the number of available tools. For instance, Gorilla (Patil et al., 2023) fine-tunes models on a massive corpus of API calls from hubs like TorchHub and TensorFlow Hub. Similarly, HuggingGPT (Shen et al., 2023) positions an LLM as a controller to delegate tasks to specialized models within the Hugging Face ecosystem. While powerful, their success hinges on environments where tools are well-documented and standardized.
Exploring Unstructured Open-Source Repositories and Challenges. A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source. Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories. These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents. They lack standardized interfaces (Zhang et al., 2024a; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task. This chaotic integration process illustrates a critical failure on the tool supply side.
The Emergence of Standardization and Unaddressed Gap. Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP. For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of 4400+ available options. This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025). This highlights a crucial, unaddressed gap: how is this ecosystem of MCP services populated in the first place?
Figure 2: Overview of the Code2MCP framework. The system takes a GitHub repository URL as input and automatically generates a complete MCP service through a multi-agent workflow.
3 Methodology: The Code2MCP Framework
To achieve the goal of automatically transforming an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents. The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request. For completeness, a more formal description of this workflow is provided in the appendix A.3.
@@ 3 Methodology: The Code2MCP Framework
Finalization and Delivery. After the core loop succeeds, the Finalize Agent organizes and packages the validated MCP service files. To facilitate review and adoption by the original repository maintainers, it also generates a README file explaining how to use the new MCP service. All artifacts are arranged into a reproducible directory structure under the workspace (Appendix A.2), and the agent can prepare a pull request to submit these additions back to the original repository.
4 Evaluation
4.1 Experimental Setup
@@
Task. The evaluation task is: given a GitHub repository URL, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure. The evaluation focuses on the overall conversion success rate across different domains and repository characteristics, and on the time and stability of Code2MCP compared with manual implementation and a GPT-4 template-based baseline on the same set of repositories.
@@
Repositories. Code2MCP is evaluated on 50 open-source repositories from 10 scientific and engineering domains (5 per domain), covering biomedical science, psychology, mathematical computing, earth science, chemistry, physics, astronomy, social science, linguistics, and econometrics. These repositories range from small utility libraries to large frameworks with complex dependency stacks. For each repository, the evaluation logs the environment construction result, MCP conversion result, number of generated tools, and failure types observed across all Run-Review-Fix rounds; per-repository details and the associated failure labels are summarized in a large table in the Appendix C, and a subset of representative scientific-computing repositories is further used for the Run-Review-Fix ablation and qualitative case studies.
Success criterion. A conversion is regarded as successful only if the following three conditions all hold: the runtime environment can be reconstructed according to the dependency specifications provided by the repository, the generated MCP server can be started and must pass a unified RunNode test, and at least three core tools can be invoked correctly in automatic tests and return valid outputs. If any of these conditions is not satisfied, the repository is counted as a failure.
Failure taxonomy. To analyze failure modes, all failed repositories are annotated with one or more labels from the following six categories:
@@
Table 2: Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
Domain Repos Env succ. Test succ. RRF recovered Avg rounds MCP succ.
@@
4.2 Large-scale Repository Conversion and Failure Analysis
Overall and per-domain success. The overall conversion success rate is first analyzed at the domain level. In Table 2, for each domain, the number of repositories with successful environment setup and basic test passing, the number of initially failing repositories recovered by the Run-Review-Fix loop and their average rounds, and the final number of successful MCP conversions. The per-domain MCP success rates range from 0.40 to 0.80: domains dominated by library-style projects, such as mathematical computing, astronomy, and linguistics, tend to have higher success rates, while domains such as physics and econometrics, which contain more complex environments and mixtures of scripts and workflow-style code, tend to have lower success rates. All statistics in the table are computed on a per-repository basis, where each repository contributes a single final outcome after the Run-Review-Fix loop. Per-repository results, including environment construction and MCP conversion outcomes as well as the number of generated tools, are summarized in the large table in the appendix and can be further grouped by domain or repository size if desired.
@@ 4.2 Large-scale Repository Conversion and Failure Analysis
Figure 3: Distribution of 42 failure labels assigned to the 18 failed repositories. A single repository can trigger multiple failure types across different Run-Review-Fix rounds.
@@
Failure modes. To understand why conversions fail on real repositories, all 18 failed repositories in the 50-repository evaluation are annotated using the six failure labels introduced in the failure taxonomy. Labels are aggregated across all Run-Review-Fix rounds, so a single repository may receive multiple labels if different error types are observed during different stages of the pipeline. The distribution of these labels is visualized in Figure 3. The labels env_failure and api_inference_error account for more than half of all failures, indicating that environment reconstruction and interface inference are the primary bottlenecks of the current workflow. The labels import_error and untoolable_repo are also relatively common, reflecting the difficulty of handling complex import paths, initialization order, and script-driven repositories in a fully automated pipeline.
@@
Run-Review-Fix ablation. To assess the specific contribution of the Run-Review-Fix self-correction loop, an ablation study is conducted on a set of representative scientific-computing repositories by comparing two configurations: a single-pass configuration without the loop and the full Code2MCP pipeline with the loop enabled. Figure 4 summarizes the results. The bar chart on the left shows that the per-repository success rates increase for most repositories once the loop is enabled, often moving from medium success under a single-generation setting to success rates close to or on par with human-written wrappers. The line chart on the right plots the average number of remaining errors as a function of the number of Run-Review-Fix rounds, where “remaining errors” denote the number of failing assertions or uncaught exceptions in the unified test after each round. These results indicate that the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main current limitation of the method.
4.3 Comparison with Human and GPT-4 Baselines

### v4

# Code2MCP: Transforming Code Repositories into MCP Services

@@ Abstract.
The Model Context Protocol (MCP) aims to create a standard for how Large Language Models use tools. However, most current research focuses on selecting tools from an existing pool. A more fundamental, yet largely overlooked, problem is how to populate this pool by converting the vast number of existing software projects into MCP-compatible services. To bridge this gap, we introduce Code2MCP, an agent-based framework that automatically transforms a GitHub repository into a functional MCP service with minimal human intervention. Code2MCP employs a multi-agent workflow for code analysis, environment setup, tool function design, and service generation, enhanced by a self-correcting loop to ensure reliability. We demonstrate that Code2MCP successfully transforms open-source computing libraries in scientific fields such as bioinformatics, mathematics, and fluid dynamics that are not available in existing MCP servers. By providing a novel automated pathway to unlock GitHub, the world’s largest code repository, for the MCP ecosystem, Code2MCP serves as a catalyst to significantly accelerate the protocol’s adoption and practical application. The code is public at https://anonymous.4open.science/r/code2mcp-7E34.
Agentic AI, Multi-Agent, Model Context Protocol, Large Language Model, GitHub
copyright: none
1. Introduction
@@
The landscape of artificial intelligence is increasingly defined by autonomous agents that leverage Large Language Models (LLMs) to interact with external tools (Wang et al., 2024). To overcome the inherent limitations of LLMs in tasks requiring real-time information or precise computation, the paradigm of tool-augmented reasoning has become central (Huang et al., 2024; Hao et al., 2023). Seminal works have demonstrated that models can effectively learn to invoke external functions (Schick et al., 2023; Qin et al., 2024b; Parisi et al., 2022). For example, Toolformer (Schick et al., 2023) shows that LLMs can be trained to invoke external APIs in a self-supervised manner, establishing the basic feasibility of tool use. Later work further scales this paradigm by orchestrating larger tool and model collections, such as Gorilla (Patil et al., 2024) and HuggingGPT (Shen et al., 2023), showing that tool use can be extended beyond a small, fixed set of predefined APIs.
@@
However, this burgeoning ecosystem faces a fundamental scalability challenge: the $N\times M$ integration problem (Liang et al., 2024; Qin et al., 2024b; Anthropic, 2023). Each of the N models or agent applications often requires a bespoke connector for each of the M tools it must access. This results in a fragmented and inefficient system where development effort is duplicated and innovation is stifled by high integration costs (Qu et al., 2025; Shen, 2024). To address this, MCP is proposed as a universal standard that specifies how agents and tools should communicate, enabling an interoperable “plug-and-play” ecosystem (Anthropic, 2023).
@@
Figure 1. While most research focuses on the consumption of tools (right side), one bottleneck is their supply (left side). Code2MCP solves the supply problem by converting the code repository into a standardized MCP-compliant tool.
Table 1. A comparative summary of related works in tool-augmented LLMs.
Type Work Core Contribution Tool Pool Tool Selection MCP
Consumer Toolformer (Schick et al., 2023) Teaching LLM to use external tools 5 predefined tools Fine-tuning $\times$
SciToolAgent (Ding et al., 2025) Domain-specific enhancement for scientific tool utilization KG of Scientific Tools (500+ tools) Retrieval on KG $\times$
HuggingGPT (Shen et al., 2023) Increase the size of tool pool huggingface.co LLM task planning $\times$
Gorilla (Patil et al., 2024) TorchHub, TensorHub (1600+ tools) Retriever-aware training $\times$
OpenAgents (Xie et al., 2023) Tool use from open-source beyond a closed pool github.com Multi-agent planning $\times$
RepoMaster (Wang and others, 2025) Rule-based deep search $\times$
RAG-MCP (Gan and Sun, 2025) RAG for tool selection from MCP MCP.so Retrieval on MCP $\checkmark$
In response to this integration challenge, the community’s efforts have evolved, inadvertently revealing a deeper, more foundational bottleneck (Yue et al., 2025a). The initial challenge is to establish the fundamental feasibility of tool use, where a limited set of tools proves the feasibility of the paradigm (Schick et al., 2023; Patil et al., 2024; Ding et al., 2025; Qin et al., 2024a). To break past the inherent scarcity of these platforms, the focus shifts to the vast landscape of open-source repositories (Wang and others, 2025; Xie et al., 2023; Wang, 2023).
This move, however, trades a scarcity problem for a chaos problem, exposing the wild non-standardization of real-world code. The MCP emerges as a direct answer to this chaos, promising a universal interface (Yu et al., 2025; Anthropic, 2023). Yet, this leads to a critical gap: existing research focuses almost exclusively on the consumption side of MCP, using services from a presumed-to-exist pool (Gan and Sun, 2025), while the foundational supply-side problem of how to populate this pool from existing software remains largely unaddressed. We identify and formalize this largely overlooked tool supply problem in the MCP ecosystem: how to automatically transform existing code repositories into reusable, registrable, and agent-consumable MCP services, rather than relying on one-off, repository-level execution. As illustrated in Figure 1, existing approaches make different trade-offs among the large, open, and standard dimensions of tool pools, while the key bottleneck of the MCP pool lies in the insufficient supply of standardized tool services. Code2MCP is designed to act as a supplier that fills this gap.
@@
However, while these efforts advance the consumption side of the problem by improving how agents use tools, they largely overlook a more fundamental bottleneck on the supply side. This bottleneck is not theoretical. Most consumer-side settings assume that a usable tool pool already exists, whether it is a curated knowledge graph as in SciToolAgent (Ding et al., 2025) or open-source repositories as explored by OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025). Yet the scale of available standardized tools remains limited. For example, RAG-MCP (Gan and Sun, 2025) performs retrieval over a pre-existing MCP server registry, with mcp.so as a representative platform, whereas the broader software ecosystem contains hundreds of millions of public GitHub repositories. The critical question of how to create a large and diverse pool of these standardized, agent-ready tools has been left unaddressed. This creates a major adoption gap, effectively locking away the largest software repository, GitHub, from this emerging ecosystem.
Figure 2. Overview of the Code2MCP framework. The system takes a GitHub repository URL as input and automatically generates a complete MCP service through a multi-agent workflow.
@@
In this paper, we introduce Code2MCP, a new framework designed to bridge the critical tool supply gap. Code2MCP presents a blueprint for transforming a wide range of GitHub repositories into functional and documented MCP services with minimal human intervention. However, this transformation is a complex endeavor encompassing four pivotal challenges: (1) deep code comprehension to identify core functionalities, (2) reliable environment replication to ensure executability, (3) intelligent tool abstraction to design useful and valid service interfaces, and (4) robust self-correction to handle the inevitable errors throughout the process. To systematically address these challenges, as shown in Figure 2, Code2MCP implements a collaborative multi-agent system (Park et al., 2023; Yue et al., 2024; Tran et al., 2025; Guo et al., 2024; Dorri et al., 2018). Unlike general-purpose coding agents (Cognition, 2024; Jimenez et al., 2024; Zhang et al., 2024), different agents in our framework are specialized for the distinct stages of code analysis, environment setup, and API design. Crucially, the overall reliability of this workflow is ensured by an integrated Run-Review-Fix self-correction cycle, where failure traces are explicitly analyzed to generate targeted modification plans that constrain subsequent regeneration to specific files and code regions. This design differs from naive retry-based generation by enforcing localized, diagnosis-driven repairs rather than global re-synthesis, while keeping previously validated tool interfaces stable whenever possible. The key contributions of this work are listed as follows:
@@
2. Related Work
@@
Scaling Tool Availability via Structured Platforms. To overcome the limitation of fixed toolsets, subsequent research turns to large, curated platforms. These approaches significantly expand the number of available tools. For instance, Gorilla (Patil et al., 2024) fine-tunes models on a massive corpus of API calls from hubs like TorchHub and TensorFlow Hub. Similarly, HuggingGPT (Shen et al., 2023) positions an LLM as a controller to delegate tasks to specialized models within the Hugging Face ecosystem. While powerful, their success hinges on environments where tools are well-documented and standardized.
Exploring Unstructured Open-Source Repositories and Challenges. A more ambitious paradigm shift involves treating the entirety of open-source code repositories as a virtually infinite tool source. Frameworks like OpenAgents (Xie et al., 2023) and RepoMaster (Wang and others, 2025) empower agents to directly parse, reason about, and execute code within GitHub repositories. These works confront the complexity of real-world code but expose the core bottleneck: the vast majority of these repositories are not designed for programmatic use by LLM agents. They lack standardized interfaces (Zhang et al., 2025; Jin et al., 2024; Ray, 2025), forcing the agent into an ad-hoc, brittle, and unreliable process of reverse-engineering the code, setting up its environment, and managing dependencies for every single task. While these systems partially touch the supply side by enabling agents to operate over open-source repositories, they do not aim to produce standardized, reusable MCP services that can be shared across agents and tasks.
The Emergence of Standardization and Unaddressed Gap. Recognizing this chaos, the community has moved towards standardization, exemplified by the MCP. For example, RAG-MCP (Gan and Sun, 2025) explores how an agent can effectively retrieve and select the most appropriate MCP service from a pool of available options. This approach is promising, but it presumes the existence of a rich ecosystem of MCP-compliant services (Hasan et al., 2025). This highlights a crucial gap: how is this ecosystem of MCP services populated in the first place?
3. Methodology: The Code2MCP Framework
To automatically transform an arbitrary GitHub repository into a fully functional and reliable MCP service, we design Code2MCP, an automated framework driven by the collaboration of seven specialized agents. The entire conversion process, as depicted in Figure 2, is a multi-agent workflow that begins with code analysis, proceeds through a core Run-Review-Fix self-correction loop, and culminates in the generation of a merge-ready pull request. A detailed description of the framework is provided in Appendix 1.
@@ 3. Methodology: The Code2MCP Framework
Figure 3. The complete output directory structure generated by the Code2MCP framework. The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
@@
Finalization and Delivery. After the core loop succeeds, the Finalize Agent organizes and packages the validated MCP service files. To facilitate review and adoption by the original repository maintainers, it also generates a README file explaining how to use the new MCP service. As shown in Figure 3, All artifacts are arranged into a reproducible directory structure under the workspace, and the agent can prepare a pull request to submit these additions back to the original repository.
Threat model and safety boundaries. Code2MCP executes third-party code during environment setup and validation, which poses inherent security risks. In our evaluation, all conversions are run inside isolated worker environments with restricted filesystem scope and bounded runtime budgets. By default, outbound network access is disabled during validation; for repositories whose core functionality intrinsically depends on external services, network access is selectively enabled under an explicit allowlist and treated as a separate evaluation condition. We do not attempt to systematically evaluate adversarial repositories or malicious payloads in this work; instead, our security discussion focuses on operational safeguards during automated conversion rather than adversarial robustness.
4. Evaluation
4.1. Experimental Setup
@@
Task. The evaluation task is: given a GitHub repository, Code2MCP automatically generates an MCP service for that repository, and the outcome of the conversion is judged as a success or a failure. The evaluation focuses on the overall conversion success rate across different domains and repository characteristics, and on the time and stability of Code2MCP compared with manual implementation and a GPT-4 template-based baseline on the same set of repositories.
@@
Repositories. Code2MCP is evaluated on 50 open-source repositories from 10 scientific and engineering domains (5 per domain), covering biomedical science, psychology, mathematical computing, earth science, chemistry, physics, astronomy, social science, linguistics, and econometrics. The goal of this selection is not exhaustive coverage, but to expose the framework to a diverse range of dependency structures, API styles, and execution environments commonly encountered in scientific software. For each repository, the evaluation logs the environment construction result, MCP conversion result, number of generated tools, and failure types observed across all Run-Review-Fix rounds; per-repository details and the associated failure labels are summarized in a large table in the Appendix C, and a subset of representative scientific-computing repositories is further used for the Run-Review-Fix ablation and qualitative case studies.
Success criterion. A conversion is counted as successful iff all of the following hold: (1) the runtime environment can be reconstructed from repository-provided specifications; (2) the generated MCP server starts and passes the unified RunNode validation; (3) at least three pre-identified, documentation-facing functionalities succeed under automatic invocation. These functionalities are selected deterministically from the repository’s public documentation and matched to generated tools by name and semantic description, prior to conversion.
Failure taxonomy. Each failed repository is annotated with one or more failure labels derived from execution artifacts (installation logs, server logs, and schema validation). If multiple error types occur across Run–Review–Fix rounds, labels are aggregated; a primary label is also recorded based on the final failing stage.
@@
Table 2. Per-domain summary of environment setup success, basic test success, recovery by the Run-Review-Fix (RRF) loop (“–” if none are recovered), and final MCP conversion success.
Domain Repos Env success Test success RRF recovered Avg rounds MCP success
@@
4.2. Large-scale Repository Conversion and Failure Analysis
Overall and per-domain success. The overall conversion success rate is first analyzed at the domain level.
(a) Failure label distribution.
(b) RRF ablation results.
@@ 4.2. Large-scale Repository Conversion and Failure Analysis
Figure 4. (a) summarizes the distribution of 42 failure labels assigned to 18 failed repositories; a single repository can trigger multiple failure types across different Run-Review-Fix rounds. (b) reports repository success rates with and without Run‑Review‑Fix and the remaining errors across RRF rounds for five representative scientific MCP repositories.
In Table 2, for each domain, the number of repositories with successful environment setup and basic test passing, the number of initially failing repositories recovered by the Run-Review-Fix loop and their average rounds, and the final number of successful MCP conversions. The per-domain MCP success rates range from 0.40 to 0.80: domains dominated by library-style projects, such as mathematical computing, astronomy, and linguistics, tend to have higher success rates, while domains such as physics and econometrics, which contain more complex environments and mixtures of scripts and workflow-style code, tend to have lower success rates. All statistics in the table are computed on a per-repository basis, where each repository contributes a single final outcome after the Run-Review-Fix loop. Per-repository results, including environment construction and MCP conversion outcomes as well as the number of generated tools, are summarized in the large table in the appendix and can be further grouped by domain or repository size if desired.
@@
Failure modes. To understand why conversions fail on real repositories, all 18 failed repositories in the 50-repository evaluation are annotated using the six failure labels introduced in the failure taxonomy. Labels are aggregated across all Run-Review-Fix rounds, so a single repository may receive multiple labels if different error types are observed during different stages of the pipeline. The distribution of these labels is visualized in Figure 4(a).
The labels env_failure and api_inference_error account for more than half of all failures, indicating that environment reconstruction and interface inference are the primary bottlenecks of the current workflow. The labels import_error and untoolable_repo are also relatively common, reflecting the difficulty of handling complex import paths, initialization order, and script-driven repositories in a fully automated pipeline.
Figure 5. Task success rates (top row) and completion time (bottom row) for Human experts, GPT-4 with a template, and Code2MCP across representative repositories in four repository groups. The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
@@
Run-Review-Fix ablation. To assess the specific contribution of the Run-Review-Fix self-correction loop, an ablation study is conducted on a set of representative scientific-computing repositories by comparing two configurations: a single-pass configuration without the loop and the full Code2MCP pipeline with the loop enabled. Figure 4(b) summarizes the results. The bar chart on the left shows that the per-repository success rates increase for most repositories once the loop is enabled, often moving from medium success under a single-generation setting to success rates close to or on par with human-written wrappers. The line chart on the right plots the average number of remaining errors as a function of the number of Run-Review-Fix rounds, where “remaining errors” denote the number of failing assertions or uncaught exceptions in the unified test after each round. These results indicate the majority of initial failures can be automatically repaired within one to three Run-Review-Fix rounds, whereas the few repositories that still fail after multiple rounds represent the main limitation of the method.
4.3. Comparison with Human and GPT-4 Baselines